OpenStudy (anonymous):

A company bought 1,000 shares at Rs. 56.27 per share and sold them after 1 year at Rs. 65.25. With a 1.2% commission rate of buying and selling the stock. What is its return on investment?

OpenStudy (anonymous):

pakistan say ho ?

OpenStudy (anonymous):

yeah buddy

OpenStudy (anonymous):

just wanna confirm that my own answer is right

OpenStudy (anonymous):

how come pakistan say bohat saray log hain yahan ?

OpenStudy (anonymous):

you want the answer in percentage ? or rs amount ?

OpenStudy (anonymous):

both if you have already solved it

OpenStudy (anonymous):

let me take out my calculator, im studying bio right now lol

OpenStudy (anonymous):

lol thanks mate

OpenStudy (anonymous):

medal do mujhay question is illogical. how can someone get commission on both selling and buying. its one way or another unless you are a broker, but then its not your return on investment

OpenStudy (anonymous):

yaar say the company is not a broker then what would be answer

OpenStudy (anonymous):

it can be 1.0627 or .903

OpenStudy (anonymous):

kaya jab sell karey gi tab commission de gi ya buy karey gi tab commission dey gi, i am not sure about this that's why i have got two answers

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

.10627

OpenStudy (anonymous):

see this solution shares at Rs.56.27 = 56270 Commission at 1.2% = 56270*1.2/100 = 675.24 Total cost = 56270+ 675.24 = 56945.24 Sold: 1000 shares at Rs. = 65250 Commission at 1.2% = 65250*1.2/100 = 783 Total Sale = 65250- 783 = 64467 Gain Net receipts = 64467 Total cost = 56945.24 Net Gain = 64467 – 56945.24 = 7521.76 Return on investment = 7521.76/56945.24*100 = 13.20%

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

return - investment/investment = ROI

OpenStudy (anonymous):

so my solution was right?

OpenStudy (anonymous):

bhai sara masla question ka hay, tumhara jawab bhi sahi hay ager commisiion pay kero sale per, mera jawab main main nay commision earn kia both ways so it is different. lakin method same hay

OpenStudy (anonymous):

actually tumhara formula bhi ghalat hay for ROI

OpenStudy (anonymous):

lol yeah same thing i have got

OpenStudy (anonymous):

ok leave it i have a second question

OpenStudy (anonymous):

In a domestic business an owner invests 15,000 Rs at 15% per annum for a period of 66 months. Calculate compound interest and also future value. Compound Interest Invested Amount=P=15,000 Interest Rate= I= 15% Period= 66/12=5.5 years S = P(1 + r/100)^ n S= 15000(1+15/100) ^5.5 S(Which is also the future value) =32354.1 Rs Now Compound Interest is Compound interest = S – P Compound interest=32354.1-15000 Compound interest=17354.1 Rs Future Value=32354.1 Rs we can also find fv by this formula fv (future value)= pv(present value)*(1+i)^n

OpenStudy (anonymous):

solution is with the post

OpenStudy (anonymous):

so whats the issue ?

OpenStudy (anonymous):

mean is that right what i have done?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok one last question with solution

OpenStudy (anonymous):

How much money deposited now will provide payments of Rs. 10,000 at the end of each quarter for 15 years if interest is 12% compounded quarterly. Amount of annuity (R) = 10,000 Rs Rate of interest (I) = 12% / 4 = 0.030 Number of periods = n = 15 ×4 = 60 PV =R* [(1-(1+i) ^- n)/I] PV = 10000*[1-(1+0.12)^-60/0.12] PV = 10000*[0.9907] PV= 9907.16 RS

OpenStudy (anonymous):

my formula and solution is right?

OpenStudy (anonymous):

number of periods is wrong

OpenStudy (anonymous):

nvm its corrent lol

OpenStudy (anonymous):

everything is good

OpenStudy (anonymous):

thanks buddy

OpenStudy (anonymous):

mujhey bhi medal de hi do

OpenStudy (anonymous):

hunk_jay@hotmail.com if you wana ask any questions, add me on facebook since we cant make friends here

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

adding you

OpenStudy (anonymous):

k

OpenStudy (anonymous):

added

OpenStudy (anonymous):

k see you later

OpenStudy (anonymous):

thanks brother have fun take care bye bye

OpenStudy (anonymous):

bye